OpenStudy (anonymous):

in how many orders can 5 girls and 3 boys walk through a doorway when order does matter? when order doesn't matter?

OpenStudy (anonymous):

if order does does matter, then by the counting principle there are \[8 !=8\times 7\times 6\times 5\times 4\times 3\times 2\] ways

OpenStudy (anonymous):

the wording of the question doesn't make any sense really "n how many ORDERS can 5 girls and 3 boys walk through a doorway when ORDER does matter?"

OpenStudy (anonymous):

the second part really makes no sense what does "how many ways can 8 people walk through a door when order doesn't matter? " what on earth is that supposed to mean? i guess the answer is one, they all walk through together

OpenStudy (anonymous):

How many orders can they walk in if there are restrictions? How many orders can they walk in without restrictions? ....sorry

OpenStudy (anonymous):

what restrictions?

OpenStudy (anonymous):

girls walk through before boys

OpenStudy (anonymous):

either i am missing something or the question is illiterate

OpenStudy (anonymous):

I didn't read thoroughly enough as I typed. How many ways can the kids walk through the doorway if there are no rules/restrictions as far as who goes when? How many ways can the kids walk through if the girls must walk through before the boys

OpenStudy (anonymous):

if there are no restrictions then the answer is the one i wrote above 8 choices for the first persone 7 for the second 6 for the third 5 for the fourth 4 for the fifth 3 for the sixth 2 for the seventh 1 for the eighth by the counting principle the total number of ways to do this is \[8!=8\times 7\times 6\times 5\times 4\times 3\times 2\] i.e. you multiply the number of ways

OpenStudy (anonymous):

second one is similar, but since girls go before boys and there are 5 girls and 3 boys it is \[5!\times 3!\]

OpenStudy (anonymous):

thank you!